OpenStudy (ilovehorses14):

Which pair of numbers is relatively prime? A. 143 and 63 B. 43 and 172 C. 22 and 143 D. 110 and 77

OpenStudy (mathmale):

I'd strongly suggest that y ou look up the mathematical definition of "relatively prime," so that you 'd know precisely what y ou're looking for.

zepdrix (zepdrix):

So you're trying to determine which pair of numbers `share no factors`. How bout option D? Do you know how to list some factors? 77 = 11*7 How bout 110?

OpenStudy (ilovehorses14):

Which of the following is a composite number? A. 23 B. 11 C. 25 D. 19

OpenStudy (ilovehorses14):

A composite number is a positive integer that has at least one positive divisor other than one or the number itself. Any integer greater than one that is not a prime number is a composite number, while the number one is a unit; it is neither prime nor composite.

OpenStudy (ilovehorses14):

i dont have a book so plz help with the question i just poted

OpenStudy (ilovehorses14):

posted i meant

zepdrix (zepdrix):

25 ends in a 5, is there any way to factor this number into something * something? How bout 23? any ideas? :d

zepdrix (zepdrix):

11 = 11*1, clearly 11 is not a composite number. There are no other factors of 11.

OpenStudy (ilovehorses14):

i looked up wat the factors are for 23 and it said 1 and 23

OpenStudy (ilovehorses14):

The factors of 25 Answer : 1,5,25,

OpenStudy (ilovehorses14):

ok so 25 11 and 23 are prime so i guess the only one that is not prime is 19

OpenStudy (ilovehorses14):

am i right cuz i looked them up

OpenStudy (ilovehorses14):

nvm 19 is a prime number

OpenStudy (ilovehorses14):

im going to go with 25

OpenStudy (ilovehorses14):

Which pair of numbers is relatively prime? A. 5 and 75 B. 12 and 51 C. 7 and 20 D. 15 and 21

zepdrix (zepdrix):

Yes, 25 has 5 as a factor. So it is a composite number :)

zepdrix (zepdrix):

5 and 75 both end in 5, so they share a factor of 5. They are not relatively prime.

OpenStudy (ilovehorses14):

so a is not right

zepdrix (zepdrix):

A bunch of the options have 3 as one of their factors. Here is an easy place to start maybe. Start with the one that you know is a prime number. 7 is a prime number. So 7 and 20 are relatively prime unless 20 has a factor of 7. Does 20 = 7*(something)?

OpenStudy (ilovehorses14):

Which pair of numbers is relatively prime? A. 108 and 57 B. 21 and 183 C. 55 and 121 D. 60 and 151

OpenStudy (ilovehorses14):

anyone know the answer

OpenStudy (ilovehorses14):

q .5 Which pair of numbers is relatively prime? A. 102 and 321 B. 10 and 45 C. 3 and 51 D. 35 and 72